# Jet transport coefficients by elastic and radiative scatterings in the strongly interacting quark-gluon plasma

### v1

@@ V.1 Dependence of transport coefficients $\hat{q}$ and $dE/dx$ on the mass of the emitted gluon
Figure 4: Temperature dependence of $\hat{q}/T^{3}$ (left plots) and of the energy loss $(1/T^{2})dE/dx$ (right plots) for elastic (blue lines) and inelastic (orange lines) scattering for different momenta of the jet parton: $p=10$ GeV$/c$ (upper plots) and 100 GeV$/c$ (lower plots). The inelastic cross sections for the emission of a massive gluon are shown by solid lines while for a massless ($M_{g}=0$) gluon by dashed lines.
Figure 4: Temperature dependence of $\hat{q}/T^{3}$ (left plots) and of the energy loss $(1/T^{2})dE/dx$ (right plots) for elastic (blue lines) and inelastic (orange lines) scattering for different momenta of the jet parton: $p=10$ GeV$/c$ (upper plots) and 100 GeV$/c$ (lower plots). The inelastic cross sections for the emission of a massive gluon are shown by solid lines while for a massless ($M_{g}=0$) gluon by dashed lines.
@@
Fig. 5 shows the temperature dependence of the scaled $\hat{q}$ coefficient (left plots) and the scaled energy loss $dE/dx$ (right plots) for elastic (blue lines), inelastic scattering (orange lines) with radiation of a massive gluon (solid lines) and massless ($M_{g}=0$) gluon (dashed lines) for two quark jet momenta: $p=10$ GeV$/c$ (upper plots) and $p=100$ GeV$/c$ (lower plots). Similar to differential cross sections, the $\hat{q}$ and $dE/dx$ strongly increase as the gluon mass decreases. Moreover, for the massless emitted gluon the inelastic $\hat{q}$ and $dE/dx$ become significant even for small jet momenta of 10 GeV$/c$.
Fig. 5 displays the inelastic $\hat{q}/T^{3}$ (left plot) and energy loss $(1/T)^{2}dE/dx$ (right plot) as functions of the emitted gluon mass $M_{g}$. For all displayed temperatures and jet momenta the values of $\hat{q}$ and energy loss $dE/dx$ monotonically decrease with increasing of the emitted gluon mass.
@@ V.2 Comparison of jet transport coefficients with different scenarios for the effective coupling constant
Here we investigate the influence of the effective coupling $g$ in thermal, jet, and radiative vertices on the temperature and energy dependence of $\hat{q}$ and $dE/dx$ within the different Scenarios 0-IV as described in Section III.2. We stress that all calculations of Feynman diagrams are presented within the ”standard” DQPM propagators and masses of quasiparticles taken at the pole mass of the spectral function, i.e. only the coupling constants in the vertices are varied – cf. Table 1 and Fig. 2.
@@
$\bullet$ The upper rows in Figs. 6 and 7 show the standard DQPM results (”Scenario 0”) where for all vertices the thermal $g(T)$ is used. As seen from these plots, accounting for the inelastic scattering substantially increases $\hat{q}/T^{3}$ and $(1/T^{2})dE/dx$, especially at low temperatures due to the strong $T$-dependence of the strong coupling $g(T)$ – cf. Fig. 2.
@@
$\bullet$ The second row in Figs. 6 and 7 show the results for the ”Scenario I” where all vertices (thermal, jet parton, emitted gluon) have been taken as a constant $g=\sqrt{4\pi\cdot 0.3}$, i.e. $\alpha_{s}=0.3$. This scenario allows to illustrate explicitly the influence of the choice of the coupling $g$ on $\hat{q}$ and $dE/dx$. One can see that the scaled $\hat{q}$ decreases with decreasing $T$, so the strong rise of $\hat{q}$ for ”Scenario 0” at low $T$ comes from the increase of the DQPM strong coupling $g(T)$ in the vicinity of $T_{c}$ as illustrated in Fig. 2. The same holds for the energy loss $dE/dx$, however, the relative contribution of inelastic reactions is larger than that for $\hat{q}$ and the $p$-dependence is stronger.
@@
$\bullet$ The third row in Figs. 6 and 7 shows the ”Scenario II” where the coupling constants in the jet parton and emitted gluon vertices are taken from the Zakharov model Zakharov (2021, 2020) and depend on the transverse momenta squared in the corresponding vertex – $g(Q^{2})$ or $g(k_{t}^{2})$. The scaled $\hat{q}$ for this scenario shows a small rise with decreasing $T$. The dashed lines show the dependence of $\hat{q}$ on the choice of the parameter $\alpha_{s}^{fr}$ in Eq. (12): the solid lines stand for $\alpha_{s}^{fr}=1.05$ (vacuum value) while the dashed line indicates the results for $\alpha_{s}^{fr}=0.42$ (in medium value). One can see that the in-medium modification of $\alpha_{s}^{fr}$ only slightly reduces $\hat{q}$. This reduction is more prominent for inelastic reactions depicted by the green lines. Again, similar observations are valid for $dE/dx$ with stronger $T$- and $p$- dependencies.
@@
$\bullet$ The fourth row in Figs. 6 and 7 represent the ”Scenario III” with an energy dependent $g(E)$ – in line with the QLBT model Liu et al. (2021, 2024) – in the jet and radiative vertices. This leads to a quite flat $\hat{q}$ of smaller value as well as only a small increase of $\hat{q}$ with jet energy (cf. left and right plots) compared to the other scenarios due to the fact the $g(E)$ is smaller than $g$ from the other scenarios; it doesn’t depend on $T$ and decrease with $E$ as demonstrated in Fig. 2 by the green lines. For $dE/dx$ we observe a similar trend, however, the relative contribution of inelastic reactions is smaller than for the other scenarios.
@@
$\bullet$ The fifth row in Figs. 6 and 7 displays the ”Scenario IV” in line with the couplings used in the DREENA model Zigic et al. (2022); Karmakar et al. (2023); Djordjevic and Djordjevic (2014); Djordjevic (2009) implying the HTL interaction of the jet quark with the thermal QCD medium. Here the jet vertex is replaced by the $g(ET)$ coupling and the radiative vertex by $g(Q^{2})$ where $Q^{2}$ stands for the virtuality of the intermediate parton before(after) the gluon emission. In this scenario the behaviour of $\hat{q}$ and $dE/dx$ is similar to ”Scenario III” and dominated by the energy dependence of the coupling $g(ET)$ – cf. orange lines in Fig. 2.
@@
As follows from Fig. 8 the behaviour of the scaled $\hat{q}$ is dominated by the choice of the strong coupling $g$ and deviates very strongly especially at low temperatures $T$. While ”Scenario I” with a fixed $\alpha_{s}=0.3$ shows a decrease of $\hat{q}$ at low $T$, the DQPM (”Scenario 0”) shows a strong rise due to the thermal $g(T)$ for all vertices.
@@
We stress that the difference between Scenario 0 and Scenario I is related only to the ratio $(g(T)/\sqrt{4\pi\cdot 0.3})^{6}$ which is changing the $T$-dependence of $\hat{q}(T)$ drastically. Other Scenarios II-IV include the momentum/energy dependence of the strong coupling in jet and radiative vertices that modifies $T$- as well as $E-$ dependencies of $\hat{q}$ and suppresses the low $T$-rise of Scenario 0: ”thermal” DQPM. We note that ”Scenario II” (motivated by the QLBT model) can not be associated directly with the LBT calculations He et al. (2015) since ”Scenario II” is based on the DQPM model for computing $\hat{q}$ where only jet and radiative vertices have been replaced in line with the QLBT model. We recall that a similar finding of the strong dependence of the drag coefficient $A(T)$ of a heavy quark on the strong coupling $g$ and the parton masses has been pointed out in Ref. Berrehrah et al. (2017).
@@
As seen from Fig. 8 the spread of the different model results for $\hat{q}$ is very large and strongly depends on the model assumptions as well on the method used for the extraction of $\hat{q}$ from the heavy-ion data on the ratio $R_{AA}$ (the nuclear modification factor of high $p_{T}$ transverse spectra of hadrons in $A+A$ collisions versus $p+p$ collisions scaled with $N_{\text{coll}}$ – the number of binary collisions in $A+A$) and flow coefficients $v_{n}$.
@@
Figure 9: Momentum dependence of $\hat{q}/T^{3}$ for different scenarios for a jet parton with medium temperature $T=0.2$ GeV (left side) and $T=0.4$ GeV (right side). Blue lines represent results for elastic processes only and orange lines represent results for the sum of elastic and inelastic contributions.
@@
Figure 10: Momentum dependence of the energy loss $dE/dx$ for different scenarios for a jet parton with medium temperature $T=0.2$ GeV (left side) and $T=0.4$ GeV (right side). Blue lines represent results for elastic processes only and orange lines represent results for the sum of elastic and inelastic contributions.
@@
In Figs. 9 and 10 we present the jet momentum dependence of $\hat{q}/T^{3}$ and $(1/T^{2})dE/dx$, respectively, for different scenarios for a jet parton with a medium temperature $T=0.2$ GeV (left side) and $T=0.4$ GeV (right side). The blue lines represent results for elastic processes only and the orange lines represent the results for the sum of elastic and inelastic contributions. Here one can see that the scaled $\hat{q}$ and $dE/dx$ for elastic and elastic + inelastic collisions grow with increasing of jet momentum $p$ for all Scenarios for the coupling constants. The $p$-dependence of the elastic $\hat{q}$ and $dE/dx$ is weaker than for the inelastic one.
@@
As follows from Figs. 9 and 10 the absolute values of $\hat{q}/T^{3}$ and $dE/dx$ depend strongly on the choice of the strong coupling in the interaction vertices. The assumption of a fast thermalisation of the jet parton and emitted gluon in the sQGP leads to a larger energy loss at low $T$ and a transverse-momentum broadening (Scenario 0) compared to the other scenarios with a strong coupling that depends on the energy/momentum transfer.
@@
The different scenarios for the jet parton interaction in the sQPG can be tested in future studies with realistic calculations of $R_{AA}$ and $v_{2}$ based on the PHSD microscopic transport approach Cassing (2009); Moreau et al. (2019) by incorporating the radiative processes and accounting for the LPM effects. We note that the sensitivity of the $R_{AA}$ and $v_{2}$ to the choice of the strong coupling has been addressed within the BAMPS model by comparing the results for constant $\alpha_{s}=0.3$ with those for a temperature dependent $\alpha_{s}(T)$ Senzel (2020). This sensitivity might be larger within the DQPM based description of jet attenuation in the sQGP medium due to the additional temperature dependence of the DQPM parton masses and widths.
@@ VI Relation between $\eta/s$ and $T^{3}/\hat{q}$
Finally, we explore the relation of $\hat{q}$ to the ratio of the specific shear viscosity to entropy density $\eta/s$. It’s important to recognize that the jet quenching parameter $\hat{q}/T^{3}$ serves as a direct measure of the parton coupling strength within the medium. Furthermore, generally one can anticipate that a higher coupling strength will correspond to a lower value of $\eta/s$ Majumder et al. (2007).
@@
This approximation can be tested within more flexible quasi-particle model frameworks, where we can explore properties of the QGP beyond the weakly interacting limit and grasp understanding towards the regime described by lQCD predictions for thermodynamic observables. The violation of the limit and more rigorous description of the QGP medium is especially relevant for moderate jet energies. It is generally expected that the assertion from Ref. Müller (2021), which posits the temperature dependence of $\eta/s$ as a linearly growing function of $T^{3}/\hat{q}$, holds predominantly in the high-temperature, weak coupling regime. Conversely, for temperatures $T\leq 2T_{c}$, the ratio $T^{3}/\hat{q}$ is expected to decrease rapidly, significantly deviating from the linear trajectory described by $1.25T^{3}/\hat{q}$ (for more details, see discussion in Karmakar et al. (2023)). Thus, there are two regimes of a weakly and strongly coupled medium which lead to different jet attenuation in the medium. The considered quasi-particle description differs from our model, therefore here we aim to provide our findings employing different coupling scenarios. Moreover, we include elastic and radiative scatterings.
@@
In Fig. 11 we show the specific shear viscosity $\eta/s$ as a function of the scaled temperature $T/T_{c}$ (upper plot) and the ratio of $\eta/s$ to $T^{3}/\hat{q}$ as a function of the scaled temperature $T/T_{c}$ (lower plot). The red solid line in the upper plot shows $\eta/s$ computed by the RTA (relaxation-time approximation) approach within the DQPM model Soloveva et al. (2020) in comparison to the lQCD data (symbols) for gluodynamics ($N_{f}=0$). The blue area indicates the DQPM results (Scenario 0) for $1.25T^{3}/\hat{q}$ computed for the jet momentum range of $3-10$ GeV$/c$ (following the DREENA-A selection) which defines the width of the spread results. Here $\hat{q}$ includes the elastic and inelastic parts. The DQPM result is compared to the estimates from the DREENA model Zigic et al. (2022); Karmakar et al. (2023) (green area). As expected, the scaled results $1.25T^{3}/\hat{q}$ and $\eta/s$ agree only at large $T\approx 2.5-3.5T_{c}$ while at low $T$ the ratio $T^{3}/\hat{q}$ decreases with $T$ stronger than $\eta/s$ – similar to the DREENA model.
@@
The deviation from the weak-coupling scaling (15) is demonstrated in the lower part of Fig. 11 which shows the ratio of $(\eta/s)(T)$ to $T^{3}/\hat{q}$ as a function of the scaled temperature $T/T_{c}$. The dashed gray line corresponds to the weak interaction limit 1.25 from Ref. Müller (2021). The blue area stands for the DQPM results for Scenario 0 with thermal coupling $g(T)$ and the thermal mass of the emitted gluon $M_{q}(T)$, which shows a strong deviation from 1.25 for lower $T$. The DQPM result is compared to the estimates from the DREENA model Zigic et al. (2022); Karmakar et al. (2023) (green area). Both models show a similar behavior of $\eta/s$, however, the DQPM rises faster with $T$. On the low plot we present the ratio $(\eta/s)(T)$ to $T^{3}/\hat{q}$ for the DQPM results for different scenarios of strong coupling separately: the blue area stands for the DQPM results for Scenario 0 from the upper plot; the red area illustrates the Scenario I – the DQPM with $\alpha_{s}=0.3$ and thermal mass of the emitted gluon $M_{g}(T)$, which decreases with decreasing $T/T_{c}$; the orange area stands for Scenario II (Zakharov couplings), which shows a much flatter distribution.
@@
Thus, our results demonstrate that the ratio $\eta/s$ to $T^{3}/\hat{q}$ has a strong dependence – especially when approaching $T_{c}$ – on the choice of the strong coupling $g$ in the scattering vertices. For the models with a temperature-dependent $g(T)$ as in the DQPM, the ratio deviates from a constant at low $T$ which corresponds to the strong interaction regime and it is approaching the scale of 1.25 at high $T$ as predicted in Ref. Müller (2021), where the strong coupling $g(T)$ decreases to a small value of 0.3.
@@
Figure 11: Upper plot: Specific shear viscosity $\eta/s$ obtained from Eq. (15) as a function of the scaled temperature $T/T_{c}$: the blue area corresponds to the Scenario 0 – the DQPM with thermal $g(T)$ and thermal mass of the emitted gluon $M_{q}(T)$; the green area shows the estimates from the DREENA model Karmakar et al. (2023). The widths of the areas correspond to the jet momentum range of $3-10$ GeV$/c$. The dashed gray line shows the infinitely strong coupling limit or Kovtun-Son-Starinets bound Kovtun et al. (2005) $(\eta/s)_{\text{KSS}}=1/4\pi$. The symbols display lQCD data for pure SU(3) gauge theory taken from Refs. Astrakhantsev et al. (2017) (blue circles), Nakamura and Sakai (2005) (orange circles), Sakai and Nakamura (2007) (green circles), Meyer (2007) (red circles). Lower plot: product of $\eta/s$ and $\hat{q}/T^{3}$ as a function of the scaled temperature $T/T_{c}$. The dashed gray lines show the weak interaction limit Müller (2021). The blue area corresponds to the Scenario 0 (from the upper plot), the red area shows the Scenario I – the DQPM with $\alpha_{s}=0.3$ and thermal mass of the emitted gluon $M_{g}(T)$; the orange area indicates the Scenario II (Zakharov couplings) and thermal mass of the emitted gluon $M_{g}(T)$.

### v2

@@ V.1 Dependence of transport coefficients $\hat{q}$ and $dE/dx$ on the mass of the emitted gluon
Figure 4: Temperature dependence of $\hat{q}/T^{3}$ (left plots) and the energy loss $(1/T^{2})dE/dx$ (right plots) for elastic (blue lines) and inelastic (orange lines) scattering for different momenta of the jet parton: $p=10$ GeV$/c$ (upper plots) and 100 GeV$/c$ (lower plots). The inelastic transport coefficients for the emission of a massive gluon are shown by solid lines, while for a massless ($M_{g}=0$) gluon by dashed lines.
@@
Figure 5 shows the temperature dependence of the scaled $\hat{q}$ coefficient (left plots) and the scaled energy loss $dE/dx$ (right plots) for elastic (blue lines), inelastic scattering (orange lines) with radiation of a massive gluon (solid lines) and massless ($M_{g}=0$) gluon (dashed lines) for two quark jet momenta: $p=10$ GeV$/c$ (upper plots) and $p=100$ GeV$/c$ (lower plots). Similar to differential cross sections, the $\hat{q}$ and $dE/dx$ increase as the gluon mass decreases. Moreover, for the massless emitted gluon the inelastic $\hat{q}$ and $dE/dx$ become significant even for small jet momenta of 10 GeV$/c$.
As follows from Fig. 5, the inelastic reactions contribute dominantly to $\hat{q}$ and $dE/dx$. As mentioned above, in a thermalized QGP the contribution of inelastic reactions to the interaction rate is small due to predominantly low $\sqrt{s}$ collisions of thermal partons. However, this changes for the scattering of a fast jet quark with a thermal parton at large $\sqrt{s}$. Since the inelastic $q+q$ and $q+g$ cross sections grow faster with increasing $\sqrt{s}$ than the elastic ones, the inelastic energy loss and transverse momentum broadening dominate the elastic ones for high momentum jet partons propagating through the thermalized sQGP. The available collision energy is sufficient for the emission of massless pQCD gluons (dashed lines) as well as for the heavy DQPM gluons (solid lines).
Figure 5 displays the inelastic $\hat{q}/T^{3}$ coefficient (left plot) and energy loss $(1/T)^{2}dE/dx$ (right plot) as functions of the emitted gluon mass $M_{g}$. For all displayed temperatures and jet momenta the values of $\hat{q}$ and energy loss $dE/dx$ monotonically decrease with increasing of the emitted gluon mass.
@@ V.2 Comparison of jet transport coefficients with different scenarios for the effective coupling constant
Here we investigate the influence of the effective coupling $g$ in thermal, jet, and radiative vertices on the temperature and energy dependence of $\hat{q}$ and $dE/dx$ within the different Scenarios 0-IV as described in Sec. III.2. We stress that all calculations of Feynman diagrams are presented within the ”standard” DQPM propagators, and masses of quasiparticles are taken at the pole mass of the spectral function, i.e., only the coupling constants in the vertices are varied – cf. Table 1 and Fig. 2.
@@
$\bullet$ The upper rows in Figs. 6 and 7 show the standard DQPM results (Scenario 0), where for all vertices the thermal $g(T)$ is used. As seen from these plots, accounting for the inelastic scattering substantially increases $\hat{q}/T^{3}$ and $(1/T^{2})dE/dx$, especially at low temperatures due to the strong $T$ dependence of the strong coupling $g(T)$ – cf. Fig. 2.
@@
$\bullet$ The second row in Figs. 6 and 7 show the results for the Scenario I, where all vertices (thermal, jet parton, emitted gluon) have been taken as a constant $g=\sqrt{4\pi\cdot 0.3}$, i.e., $\alpha_{s}=0.3$. This scenario allows to illustrate explicitly the influence of the choice of the coupling $g$ on $\hat{q}$ and $dE/dx$. One can see that the scaled $\hat{q}$ decreases with decreasing $T$, so the strong rise of $\hat{q}$ for Scenario 0 at low $T$ comes from the increase of the DQPM strong coupling $g(T)$ in the vicinity of $T_{c}$, as illustrated in Fig. 2. The same holds for the energy loss $dE/dx$, however, the relative contribution of inelastic reactions is larger than that for $\hat{q}$, and the $p$ dependence is stronger.
@@
$\bullet$ The third row in Figs. 6 and 7 shows the Scenario II, where the coupling constants in the jet parton and emitted gluon vertices are taken from the Zakharov model Zakharov (2021, 2020) and depend on the squared momenta in the corresponding vertex – $g(Q^{2})$ or $g(k_{T}^{2})$. The scaled $\hat{q}$ for this scenario shows a small rise with decreasing $T$. The dashed lines show the dependence of $\hat{q}$ on the choice of the parameter $\alpha_{s}^{fr}$ in Eq. (12): the solid lines stand for $\alpha_{s}^{fr}=1.05$ (vacuum value), while the dashed line indicates the results for $\alpha_{s}^{fr}=0.42$ (in-medium value). One can see that the in-medium modification of $\alpha_{s}^{fr}$ only slightly reduces $\hat{q}$. This reduction is more prominent for inelastic reactions depicted by the green lines. Again, similar observations are valid for $dE/dx$ with stronger $T$ and $p$ dependencies.
@@
$\bullet$ The fourth row in Figs. 6 and 7 represent the Scenario III with an energy dependent $g(E)$ – in line with the QLBT model Liu et al. (2022, 2024) – in the jet and radiative vertices. This leads to a quite flat $\hat{q}$ of smaller value as well as only a small increase of $\hat{q}$ with jet energy (cf. left and right plots) compared to the other scenarios due to the fact the $g(E)$ is smaller than $g$ from the other scenarios; it does not depend on $T$ and decrease with $E$ as demonstrated in Fig. 2 by the green lines. For $dE/dx$ we observe a similar trend, however, the relative contribution of inelastic reactions is smaller than for the other scenarios.
@@
$\bullet$ The fifth row in Figs. 6 and 7 displays the Scenario IV with the couplings used in the DREENA model Zigic et al. (2022); Karmakar et al. (2023); Djordjevic and Djordjevic (2014); Djordjevic (2009), implying the HTL interaction of the jet quark with the thermal QCD medium. Here the jet vertex is replaced by the $g(ET)$ coupling, and the radiative vertex by $g(Q^{2})$, where $Q^{2}$ stands for the virtuality of the intermediate parton before(after) the gluon emission. In this scenario the behavior of $\hat{q}$ and $dE/dx$ is similar to Scenario III and dominated by the energy dependence of the coupling $g(ET)$ – cf. orange lines in Fig. 2.
@@
As follows from Fig. 8, the behavior of the scaled $\hat{q}$ is dominated by the choice of the strong coupling $g$ and deviates very strongly, especially at low temperatures $T$. While Scenario I with a fixed $\alpha_{s}=0.3$ shows a decrease of $\hat{q}$ at low $T$, the DQPM (Scenario 0) shows a strong rise due to the thermal $g(T)$ for all vertices.
@@
We stress that the difference between Scenario 0 and Scenario I is related only to the ratio $(g(T)/\sqrt{4\pi\cdot 0.3})^{6}$, which changes the $T$ dependence of $\hat{q}(T)$ drastically. Other Scenarios II-IV include the momentum/energy dependence of the strong coupling in jet and radiative vertices, which modifies $T$ as well as $E$ dependencies of $\hat{q}$ and suppresses the low $T$-rise of Scenario 0. We note that Scenario II (motivated by the QLBT model) can not be associated directly with the LBT calculations He et al. (2015), since it is based on the DQPM, where only jet and radiative vertices have been replaced in line with the QLBT model. We recall that a similar finding of the strong dependence of the drag coefficient $A(T)$ of a heavy quark on the strong coupling $g$ and the parton masses has been pointed out in Ref. Berrehrah et al. (2017).
@@
As seen from Fig. 8, the spread of the different model results for $\hat{q}$ is very large and strongly depends on the model assumptions as well on the method used for the extraction of $\hat{q}$ from the heavy-ion data on the ratio $R_{AA}$ (the nuclear modification factor of high $p_{T}$ transverse spectra of hadrons in $A+A$ collisions versus $p+p$ collisions scaled with $N_{\text{coll}}$ – the number of binary collisions in $A+A$) and flow coefficients $v_{n}$.
@@
Figure 9: Momentum dependence of the scaled $\hat{q}$ coefficient for different scenarios for a jet parton with medium temperature $T=0.2$ GeV (left side) and $T=0.4$ GeV (right side). Blue lines represent results for elastic processes only and orange lines represent results for the sum of elastic and inelastic contributions.
@@
Figure 10: Momentum dependence of the scaled energy loss $dE/dx$ for different scenarios for a jet parton with medium temperature $T=0.2$ GeV (left side) and $T=0.4$ GeV (right side). Blue lines represent results for elastic processes only and orange lines represent results for the sum of elastic and inelastic contributions.
@@
In Figs. 9 and 10 we present the jet momentum dependence of $\hat{q}/T^{3}$ and $(1/T^{2})dE/dx$, respectively, for different scenarios for a jet parton at medium temperature $T=0.2$ GeV (left side) and $T=0.4$ GeV (right side). The blue lines represent results for elastic processes only, and the orange lines represent the results for the sum of elastic and inelastic contributions. Here one can see that the scaled $\hat{q}$ and $dE/dx$ for elastic and elastic + inelastic collisions grow with increasing of jet momentum $p$ for all scenarios for the coupling constants. The $p$ dependence of the elastic $\hat{q}$ and $dE/dx$ is weaker than for the inelastic one.
@@
As follows from Figs. 9 and 10, the absolute values of $\hat{q}/T^{3}$ and $dE/dx$ depend strongly on the choice of the strong coupling in the interaction vertices. The default DQPM (Scenario 0) leads to a larger transverse momentum broadening (Fig. 9) and a larger energy loss (Fig. 10) at low $T$ compared to the other scenarios. This is due to the fact that the DQPM strong coupling $g$ grows rapidly at $T\to T_{c}$, and at low $T$ it is much larger than $g$ for the other scenarios, where its value depends on the energy/momentum transfer (cf. Fig. 2). In other words, Scenario 0 corresponds to a fast ”thermalization” of the jet parton and emitted gluon in the sQGP, since the $g$ of DQPM is evaluated for the thermal medium.
@@
The different scenarios for the jet parton interaction in the sQPG can be tested in future studies with realistic calculations of $R_{AA}$ and $v_{2}$ based on the PHSD microscopic transport approach Cassing (2009); Moreau et al. (2019) by incorporating the radiative processes and accounting for the LPM effect. We note that the sensitivity of the $R_{AA}$ and $v_{2}$ to the choice of the strong coupling has been addressed within the BAMPS model by comparing the results for constant $\alpha_{s}=0.3$ with those for a temperature dependent $\alpha_{s}(T)$ Senzel (2020). This sensitivity might be larger within the DQPM based description of jet attenuation in the sQGP medium due to the additional temperature dependence of the DQPM parton masses and widths.
@@ VI Relation between $\eta/s$ and $T^{3}/\hat{q}$
Finally, we explore the relation of elastic $\hat{q}$ to the ratio of the specific shear viscosity to entropy density $\eta/s$. It is important to recognize that the jet quenching parameter $\hat{q}/T^{3}$ serves as a direct measure of the parton coupling strength within the medium. Furthermore, generally one can anticipate that a higher coupling strength will correspond to a lower value of $\eta/s$ Majumder et al. (2007).
@@
This approximation can be tested within more flexible quasiparticle model frameworks, where we can explore properties of the QGP beyond the weakly interacting limit and grasp understanding towards the regime described by lQCD predictions for thermodynamic observables. The violation of the limit and more rigorous description of the QGP medium is especially relevant for moderate jet energies. It is generally expected that the assertion from Ref. Müller (2021), which posits the temperature dependence of $\eta/s$ as a linearly growing function of $T^{3}/\hat{q}$, holds predominantly in the high-temperature, weak coupling regime. Conversely, for temperatures $T\leq 2T_{c}$, the ratio $T^{3}/\hat{q}$ is expected to decrease rapidly, significantly deviating from the linear trajectory described by $1.25T^{3}/\hat{q}$ (for more details, see discussion in Karmakar et al. (2023)). Thus, there are two regimes of a weakly and strongly coupled medium which lead to different jet attenuation in the medium. The considered quasiparticle description differs from our model, and therefore, here we aim to provide our findings employing different coupling scenarios for the thermal and jet vertices for the elastic scattering.
@@
In Fig. 11 we show the specific shear viscosity $\eta/s$ as a function of the scaled temperature $T/T_{c}$ (upper plot) and the ratio of $\eta/s$ to $T^{3}/\hat{q}$ as a function of the scaled temperature $T/T_{c}$ (lower plot). The red solid line in the upper plot shows $\eta/s$ computed by the RTA (relaxation-time approximation) approach within the DQPM Soloveva et al. (2020) in comparison to the lQCD data (symbols) for gluodynamics ($N_{f}=0$). The blue area indicates the DQPM results (Scenario 0) for $1.25\,T^{3}/\hat{q}$ computed for the jet momentum range of $3-10$ GeV$/c$ (following the DREENA-A selection), which defines the width of the spread results. The DQPM result is compared to the estimates from the DREENA model Zigic et al. (2022); Karmakar et al. (2023) (green area). As expected, the scaled results $1.25\,T^{3}/\hat{q}$ and $\eta/s$ agree only at large $T\approx 2-3.5T_{c}$, while at low $T$ the ratio $T^{3}/\hat{q}$ decreases with $T$ stronger than $\eta/s$ – similar to the DREENA model.
@@
The deviation from the weak-coupling scaling (15) is demonstrated in the lower part of Fig. 11, which shows the ratio of $(\eta/s)(T)$ to $T^{3}/\hat{q}$ as a function of the scaled temperature $T/T_{c}$. The dashed gray line corresponds to the weak interaction limit 1.25 from Ref. Müller (2021). The blue area stands for the DQPM results for Scenario 0, with thermal coupling $g(T)$ for the thermal and jet vertices, which shows a strong deviation from 1.25 for lower $T$. The red area illustrates the Scenario I – the DQPM with $\alpha_{s}=0.3$; the orange area stands for Scenario II (Zakharov coupling in jet vertex). These scenarios show much flatter distributions.
@@
Thus, our results demonstrate that the ratio $\eta/s$ to $T^{3}/\hat{q}$ has a strong dependence, especially when approaching $T_{c}$, on the choice of the strong coupling $g$ in the scattering vertices. For the models with a temperature-dependent $g(T)$, as in the DQPM, the ratio deviates from a constant at low $T$, which corresponds to the strong interaction regime, and it approaches the scale of 1.25 at high $T$ as predicted in Ref. Müller (2021), where the strong coupling $g(T)$ decreases to a small value of 0.3.
@@
Figure 11: Upper plot: specific shear viscosity $\eta/s$ obtained from Eq. (15) as a function of the scaled temperature $T/T_{c}$: the blue area corresponds to the Scenario 0 – the DQPM with thermal $g(T)$; the red solid line shows $\eta/s$ computed by the RTA approach within the DQPM from Ref. Soloveva et al. (2020); the green area shows the estimates from the DREENA model Karmakar et al. (2023). The widths of the areas correspond to the jet momentum range of $3-10$ GeV$/c$. The dashed gray line shows the infinitely strong coupling limit or Kovtun-Son-Starinets bound Kovtun et al. (2005) $(\eta/s)_{\text{KSS}}=1/4\pi$. The symbols display lQCD data for pure SU(3) gauge theory taken from Refs. Astrakhantsev et al. (2017) (blue circles), Nakamura and Sakai (2005) (orange circles), Sakai and Nakamura (2007) (green circles), Meyer (2007) (red circles). Lower plot: product of $\eta/s$ and $\hat{q}/T^{3}$ as a function of the scaled temperature $T/T_{c}$. The dashed gray line shows the weak interaction limit Müller (2021). The blue area corresponds to the Scenario 0 (from the upper plot) – DQPM with thermal $g(T)$, the red area shows the Scenario I – the DQPM with $\alpha_{s}=0.3$; the orange area indicates the Scenario II (Zakharov coupling in jet vertex).